ps2frag(1) ps2frag(1) NAME ps2frag-obsoleteshellscriptforthePSfragsystem. IMPORTANTNOTICE ThenewPSfragsystemnolongerrequirestheps2fragscript;instead,ithandlestheprocessing entirelywithinTeX/LaTeXandDVIPS.I’msureyouwillagreethatneverneedingtorunps2frag againisaniceconvenience! However,therearetwosignificantdifferencesinthewaythisnewversionofPSfragworks.Please makeyourselfawareofthem: 1)XDviisnolongerabletodeterminewhereyourPSfragreplacementsshouldgo,soinsteadit linesthemupinaverticallisttotheleftofthefigure.Thisallowsyoutoconfirmthattheyhave beentypesetproperly,atleast.However,toconfirmthatPSfragpositionsyourreplacements properly,youwillhavetoviewthePostScriptversionofyourfilewithaviewerlikeGhostView,or printitout.Thisseemstobetheonlydisadvantagetotheeliminationofthepre-processingstep. 2)Ifyouembed’\tex’commandsinsideyourfigures,younowneedtoexplicitlytellPSfragto processthesecommands.Todoso,use\usepackage[scanall]{psfrag}insteadof\usepackage{psfrag}atthebeginningofyourLaTeXfile. Ifyouonlyuse’\tex’commandsinasmallnumberoffigures,thenamoreefficientmightbeto turnon’\tex’-scanningonlyforthosefigures.Todothat,addthecommand’\psfragscanon’ immediatelybeforeeachrelevant\includegraphicsor\epsfboxcommand. NOTES SeethePSfragdocumentationforfurtherinformation. SEEALSO dvips(1),gs(1),ghostview(1),latex(1) AUTHORS psfrag@rascals.stanford.edu ThePSfragmaintainer’smailinglist. TeXware Feb95 1